# Targeted Data Fusion for Causal Survival Analysis Under Distribution Shift

### v1

@@ D.1.2 Lemmata for the CCOD estimator
$=\mathbb{P}_{n}[\varphi^{*\text{CCOD}}_{\infty,t,a}]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}(\mathbb{P}_{n}^{m}-\mathbb{P})\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]$
$=\mathbb{P}_{n}[\varphi^{*\text{CCOD}}_{\infty,t,a}]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\theta^{0}(t,a)\right].$ (3)
@@
Under Assumption 2.3 and Condition 2.2, there exists a universal constant $C=C(\eta)$ such that for each $m$, $n$, $t$, and $a$,
@@
Note that the expression of $\widehat{\bar{B}}_{m}(\mathcal{V}_{m})-\bar{B}_{\infty}(\mathcal{V}_{m})$ can be found in Lemma 3 of Westling et al. (2023), while we only need to replace corresponding nuisance functions with the global version here, thus the detail is omitted. Then, by the triangle inequality, we have $\mathbb{P}\left[\widehat{\varphi}^{\text{CCOD}}_{t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]^{2}\leq\left[\sum_{j=1}^{4}\{\mathbb{P}\{(\bar{U}_{j,n,m,t,a})^{2}\}\}^{1/2}\right]^{2}$. Therefore, under Assumption 2.3 and Condition 2.2, there exists a universal constant $C=C(\eta)$ such that the result in the statement holds. Thus, the proof is completed. ∎
Furthermore, we investigate conditions that make the empirical process term $\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]$ to be $o_{p}(n^{-1/2})$. This requires some preliminaries from the empirical process theory (Vaart & Wellner, 2023). We first introduce the following notation and Lemma D.2. Then, we consider results for uniform convergence over $t\in[0,\tau]$.
@@
Lemma D.2 (Lemma 4 in Westling et al. (2023)).
@@
The $\mathcal{F}^{2}_{S,p^{0},q^{0},G,t_{0},a_{0}}$ is contained in the product of the classes $\{y\mapsto\mathbb{I}(y\leq t):t\in[0,t_{0}]\}$, $\{\mathbf{x}\mapsto S(t\mid a_{0},\mathbf{x}):t\in[0,t_{0}]\}$, and the singleton class $\{(\mathbf{x},a_{0},\delta,y)\mapsto\mathbb{I}(a=a_{0},\delta=1,y\leq t)q^{0}(\mathbf{x})/[p^{0}\pi(a_{0}\mid\mathbf{x})S(y\mid a_{0},\mathbf{x})G(y\mid a_{0},\mathbf{x})]\}.$ The first two classes both have covering number $\sup_{\mathbb{Q}}N(\varepsilon,\cdot,L_{2}(\mathbb{Q}))\leq 2\varepsilon^{2}$ by Lemma D.2. The third class has uniform covering number 1 for all $\varepsilon$ because it can be covered with a single ball of any positive radius. In addition, $\mathcal{F}^{2}_{S,p^{0},q^{0},G,t,a}$ is uniformly bounded by $\eta^{3}$. Therefore, based on Lemma 5.1 in Van der Laan (2006), $\sup_{\mathbb{Q}}N(\varepsilon\eta^{3},\mathcal{F}^{2}_{S,p^{0},q^{0},G,t,a},L_{2}(\mathbb{Q}))\leq 4/\varepsilon^{4}$.
Furthermore, based on Lemma 5.1 in Van der Laan (2006) again, we can rewrite $\mathcal{F}^{3}_{S,p^{0},q^{0},G,t_{0},a_{0}}=\{(\mathbf{x},a_{0},y)\mapsto\int m_{t}(u,\mathbf{x},a_{0},y)\mu^{*}(du):t\in[0,t_{0}]\}$, where
@@
Therefore, we have shown that the three classes have covering numbers bounded by $2/\varepsilon^{2}$, $4/\varepsilon^{4}$ and $4/\varepsilon^{4}$, respectively. Therefore, by Lemma 5.1 in Van der Laan (2006),
@@
If Conditions 2.12.2 hold, then $M^{-1}\sum_{m=1}^{M}n^{-1}Mn_{m}^{1/2}\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a_{0}}\right]=o_{p}(n^{-1/2})$. If Conditions 2.3 holds as well,
@@
$\frac{Mn_{m}^{1/2}}{n}\leq\frac{M(|n_{m}-n/M|+n/M)^{1/2}}{n}\leq\frac{M|n_{m}-n/M|^{1/2}+M|n/M|^{1/2}}{n}\leq\left(\frac{M}{n}\right)^{1/2}+\frac{M}{n},$
@@
$\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\sup_{u\in[0,t]}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]\right|\leq O(n^{-1/2})\frac{1}{M}\sum_{m=1}^{M}\sup_{u\in[0,t]}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]\right|,$
@@
By Theorem 2.1.14 in Van der Vaart & Wellner (1996), the covering number of $\mathcal{F}^{\text{CCOD}}_{n,m,t,a}$ is 1 for all $\varepsilon$, so the uniform entropy integral $J(1,\mathcal{F}^{\text{CCOD}}_{n,m,t,a})$ is 1 relative to the natural envelope $|\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}|$. Therefore, there is a universal constant $C^{\prime}$ such that
@@
following definition of $\bar{A}_{j,n,m,t,a}$ terms in Lemma D.8. So we showed that $M^{-1}\sum_{m=1}^{M}\mathbb{E}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]\right|$ is bounded up to $C^{\prime\prime\prime}\sum_{j=1}^{6}\mathbb{E}[\max_{m}(\bar{A}_{j,n,m,t,a})]$ for some constant $C^{\prime\prime\prime}$. It is straightforward that by Conditions D.1 and D.2, this upper bound tends to zero, which implies that $M^{-1}\sum_{m=1}^{M}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]\right|=o_{p}(1)$.
@@
$\mathbb{E}\left[\sup_{u\in[0,t]}\left|\mathbb{G}_{n}^{m}\left\{\widehat{\varphi}^{\text{CCOD}}_{n,m,u,a}-\varphi^{\text{CCOD}}_{\infty,u,a}\right\}\right|~{}\bigg{|}~{}\mathcal{T}_{m}\right]=\mathbb{E}\left[\sup_{u\in[0,t]}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,u,a}-\varphi^{\text{CCOD}}_{\infty,u,a}\right]\right|\right]=\mathbb{E}\left[\sup_{g\in\mathcal{G}_{n,m,t,a}}\left|\mathbb{G}_{n}^{m}g\right|\right],$
@@
is bounded by a constant not depending on $n$ or $m$. By Theorem 2.14.2 of Van der Vaart & Wellner (1996), there is therefore a constant $\bar{C}$ not not depending on $n$ or $m$ such that
@@
By Assumption 2.3, Conditions 2.1, 2.2, and 2.4, this bound tends to zero. ∎
@@
$\mathbb{E}\left[\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\bar{S}_{\infty}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{\bar{S}(y-\mid a,\mathbf{X})}{\bar{S}_{\infty}(y\mid a,\mathbf{X})}\left\{\frac{\bar{G}(y\mid a,\mathbf{X})\bar{\pi}(a\mid\mathbf{X})q^{0}_{\infty}(\mathbf{X})}{\bar{G}_{\infty}(y\mid a,\mathbf{X})\bar{\pi}_{\infty}(a\mid\mathbf{X})q^{0}(\mathbf{X})}-1\right\}(\bar{\Lambda}_{\infty}-\bar{\Lambda})(dy\mid a,\mathbf{X})\right].$
@@
Following a result in Lemma 1 of Westling et al. (2023), $\mathbb{E}\{H_{\bar{S}_{\infty},\bar{G}_{\infty},t,a}(Y,\Delta,\mathbf{X})\mid\mathbf{X}=\mathbf{x}\}$ equals
@@
$+\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{q^{0}_{\infty}(\mathbf{X})}{q^{0}(\mathbf{X})}\bar{S}_{\infty}(t\mid a,\mathbf{X})\frac{\bar{\pi}(a\mid\mathbf{X})}{\bar{\pi}_{\infty}(a\mid\mathbf{X})}\int_{0}^{t}\frac{\bar{S}(y-\mid a,\mathbf{X})\bar{G}(y\mid a,\mathbf{X})}{\bar{S}_{\infty}(y\mid a,\mathbf{X})\bar{G}_{\infty}(y\mid a,\mathbf{X})}(\bar{\Lambda}_{\infty}-\bar{\Lambda})(dy\mid a,\mathbf{X})\bigg{]}.$
Furthermore, by Duhamel equation in Gill & Johansen (1990), we have
@@
$\mathbb{E}\left[\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\bar{S}_{\infty}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{\bar{S}(y-\mid a,\mathbf{X})}{\bar{S}_{\infty}(y\mid a,\mathbf{X})}\left\{\frac{\bar{G}(y\mid a,\mathbf{X})\bar{\pi}(a\mid\mathbf{X})q^{0}_{\infty}(\mathbf{X})}{\bar{G}_{\infty}(y\mid a,\mathbf{X})\bar{\pi}_{\infty}(a\mid\mathbf{X})q^{0}(\mathbf{X})}-1\right\}(\bar{\Lambda}_{\infty}-\bar{\Lambda})(dy\mid a,\mathbf{X})\right].$ (4)
D.1.3 Proof of Theorem 2.9
@@
$\widehat{\theta}^{\text{CCOD}}_{n}(t,a)-\theta^{0}(t,a)$ $=\mathbb{P}_{n}[\varphi^{*\text{CCOD}}_{t,a}]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{t,a}\right]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}\left[\widehat{\varphi}^{\text{CCOD}}_{t,a}-\theta^{0}(t,a)\right].$
By Conditions 2.1 and 2.2, the second summand on the right-hand-side is $o_{p}(n^{-1/2})$ by Lemma D.4. By Lemma D.5, $\mathbb{P}[\widehat{\varphi}^{\text{CCOD}}_{t,a}]-\theta^{0}(t,a)$ equals
$\mathbb{E}\left[\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{\bar{S}(y-\mid a,\mathbf{X})}{\widehat{\bar{S}}_{m}(y\mid a,\mathbf{X})}\left\{\frac{\bar{G}(y\mid a,\mathbf{X})\bar{\pi}(a\mid\mathbf{X})\widehat{q}^{0}_{m}(\mathbf{X})}{\widehat{\bar{G}}_{m}(y\mid a,\mathbf{X})\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})q^{0}(\mathbf{X})}-1\right\}(\widehat{\bar{\Lambda}}_{m}-\bar{\Lambda})(dy\mid a,\mathbf{X})\right].$
By Duhamel equation in Gill & Johansen (1990), we have that
@@
$=\mathbb{E}\left[\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})\left\{\frac{\bar{\pi}(a\mid\mathbf{X})}{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})}-1\right\}\int_{0}^{t}\frac{\bar{S}}{\widehat{\bar{S}}_{m}}(du\mid a,\mathbf{X})\right]$
$\quad+\mathbb{E}\left[\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{\bar{\pi}(a\mid\mathbf{X})}{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})}\left\{\frac{\bar{G}(y\mid a,\mathbf{X})\widehat{q}^{0}_{m}(\mathbf{X})}{\widehat{\bar{G}}_{m}(y\mid a,\mathbf{X})q^{0}(\mathbf{X})}-1\right\}\left(\frac{\bar{S}}{\widehat{\bar{S}}_{m}}-1\right)(du\mid a,\mathbf{X})\right]$
@@
$\quad+\underbrace{\mathbb{E}\left[\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{\bar{\pi}(a\mid\mathbf{X})}{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})}\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})\int_{0}^{t}\left\{\frac{\bar{G}(y\mid a,\mathbf{X})\widehat{q}^{0}_{m}(\mathbf{X})}{\widehat{\bar{G}}_{m}(y\mid a,\mathbf{X})q^{0}(\mathbf{X})}-1\right\}\left(\frac{\bar{S}}{\widehat{\bar{S}}_{m}}-1\right)(du\mid a,\mathbf{X})\right]}_{I_{2}}.$
@@
$\quad+\underbrace{\mathbb{E}\left[\frac{\widehat{q}^{0}_{m}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{\bar{\pi}(a\mid\mathbf{X})}{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})}\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})\int_{0}^{t}\left\{\frac{\bar{G}(y\mid a,\mathbf{X})}{\widehat{\bar{G}}_{m}(y\mid a,\mathbf{X})}-1\right\}\left(\frac{\bar{S}}{\widehat{\bar{S}}_{m}}-1\right)(du\mid a,\mathbf{X})\right]}_{I_{22}}.$

### v2

@@ D.1.3 Lemmata for the CCOD estimator
$=\mathbb{P}_{n}[\varphi^{*\text{CCOD}}_{\infty,t,a}]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}(\mathbb{P}_{n}^{m}-\mathbb{P})\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]$
$\qquad+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]$
$=\mathbb{P}_{n}[\varphi^{*\text{CCOD}}_{\infty,t,a}]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]$
$\qquad+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\theta^{0}(t,a)\right].$ (4)
@@
Under Condition D.2, there exists a universal constant $C=C(\eta)$ such that for each $m$, $n$, $t$, and $a$,
@@
Note that the expression of $\widehat{\bar{B}}_{m}(\mathcal{V}_{m})-\bar{B}_{\infty}(\mathcal{V}_{m})$ can be found in Lemma 3 of Westling et al. [48], while we only need to replace corresponding nuisance functions with the global version here, thus the detail is omitted. Then, by the triangle inequality, we have $\mathbb{P}\left[\widehat{\varphi}^{\text{CCOD}}_{t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]^{2}\leq\left[\sum_{j=1}^{4}\{\mathbb{P}\{(\bar{U}_{j,n,m,t,a})^{2}\}\}^{1/2}\right]^{2}$. Therefore, under Assumption 2.3 and Condition D.2, there exists a universal constant $C=C(\eta)$ such that the result in the statement holds. Thus, the proof is completed. ∎
Furthermore, we investigate conditions that make the empirical process term $\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]$ to be $o_{p}(n^{-1/2})$. This requires some preliminaries from the empirical process theory [43]. We first introduce the following notation and Lemma D.2. Then, we consider results for uniform convergence over $t\in[0,\tau]$.
@@
Lemma D.2 (Lemma 4 in Westling et al. [48]).
@@
The $\mathcal{F}^{2}_{S,p^{0},q^{0},G,t_{0},a_{0}}$ is contained in the product of the classes $\{y\mapsto\mathbb{I}(y\leq t):t\in[0,t_{0}]\}$, $\{\mathbf{x}\mapsto S(t\mid a_{0},\mathbf{x}):t\in[0,t_{0}]\}$, and the singleton class $\{(\mathbf{x},a_{0},\delta,y)\mapsto\mathbb{I}(a=a_{0},\delta=1,y\leq t)q^{0}(\mathbf{x})/[p^{0}\pi(a_{0}\mid\mathbf{x})S(y\mid a_{0},\mathbf{x})G(y\mid a_{0},\mathbf{x})]\}.$ The first two classes both have covering number $\sup_{\mathbb{Q}}N(\varepsilon,\cdot,L_{2}(\mathbb{Q}))\leq 2\varepsilon^{2}$ by Lemma D.2. The third class has uniform covering number 1 for all $\varepsilon$ because it can be covered with a single ball of any positive radius. In addition, $\mathcal{F}^{2}_{S,p^{0},q^{0},G,t,a}$ is uniformly bounded by $\eta^{3}$. Therefore, based on Lemma 5.1 in Van der Laan [44], $\sup_{\mathbb{Q}}N(\varepsilon\eta^{3},\mathcal{F}^{2}_{S,p^{0},q^{0},G,t,a},L_{2}(\mathbb{Q}))\leq 4/\varepsilon^{4}$.
Furthermore, based on Lemma 5.1 in Van der Laan [44] again, we can rewrite $\mathcal{F}^{3}_{S,p^{0},q^{0},G,t_{0},a_{0}}=\{(\mathbf{x},a_{0},y)\mapsto\int m_{t}(u,\mathbf{x},a_{0},y)\mu^{*}(du):t\in[0,t_{0}]\}$, where
@@
Therefore, we have shown that the three classes have covering numbers bounded by $2/\varepsilon^{2}$, $4/\varepsilon^{4}$ and $4/\varepsilon^{4}$, respectively. Therefore, by Lemma 5.1 in Van der Laan [44],
@@
If Conditions D.1D.2 hold, $M^{-1}\sum_{m=1}^{M}n^{-1}Mn_{m}^{1/2}\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a_{0}}\right]=o_{p}(n^{-1/2})$. If Conditions D.4 holds as well,
@@
$\frac{Mn_{m}^{1/2}}{n}$ $\leq\frac{M(|n_{m}-n/M|+n/M)^{1/2}}{n}$
$\leq\frac{M|n_{m}-n/M|^{1/2}+M|n/M|^{1/2}}{n}$
$\leq\left(\frac{M}{n}\right)^{1/2}+\frac{M}{n},$
@@
$\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\sup_{u\in[0,t]}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]\right|$
$\leq O(n^{-1/2})\frac{1}{M}\sum_{m=1}^{M}\sup_{u\in[0,t]}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]\right|,$
@@
By Theorem 2.1.14 in van der Vaart and Wellner [46], the covering number of $\mathcal{F}^{\text{CCOD}}_{n,m,t,a}$ is 1 for all $\varepsilon$, so the uniform entropy integral $J(1,\mathcal{F}^{\text{CCOD}}_{n,m,t,a})$ is 1 relative to the natural envelope $|\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}|$. Therefore, there is a universal constant $C^{\prime}$ such that
@@
following definition of $\bar{A}_{j,n,m,t,a}$ terms in Lemma D.8, so that $M^{-1}\sum_{m=1}^{M}\mathbb{E}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]\right|$ is bounded up to $C^{\prime\prime\prime}\sum_{j=1}^{6}\mathbb{E}[\max_{m}(\bar{A}_{j,n,m,t,a})]$ for some constant $C^{\prime\prime\prime}$. It is straightforward that by Conditions D.6 and D.7, this upper bound tends to zero, so $M^{-1}\sum_{m=1}^{M}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]\right|=o_{p}(1)$.
@@
$\mathbb{E}\left[\sup_{u\in[0,t]}\left|\mathbb{G}_{n}^{m}\left\{\widehat{\varphi}^{\text{CCOD}}_{n,m,u,a}-\varphi^{\text{CCOD}}_{\infty,u,a}\right\}\right|~{}\bigg{|}~{}\mathcal{T}_{m}\right]$ $=\mathbb{E}\left[\sup_{u\in[0,t]}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,u,a}-\varphi^{\text{CCOD}}_{\infty,u,a}\right]\right|\right]$
$=\mathbb{E}\left[\sup_{g\in\mathcal{G}_{n,m,t,a}}\left|\mathbb{G}_{n}^{m}g\right|\right],$
@@
is bounded by a constant not depending on $n$ or $m$. By Theorem 2.1.14 in Van der Vaart and Wellner [46], there is therefore a constant $\bar{C}$ not not depending on $n$ or $m$ such that
@@
By Conditions D.1, D.2, and D.3, this bound tends to zero. ∎
@@
$\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\bar{S}_{\infty}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{\bar{S}(y-\mid a,\mathbf{X})}{\bar{S}_{\infty}(y\mid a,\mathbf{X})}$
×$\quad\times\left\{\frac{\bar{G}(y\mid a,\mathbf{X})\bar{\pi}(a\mid\mathbf{X})q^{0}_{\infty}(\mathbf{X})}{\bar{G}_{\infty}(y\mid a,\mathbf{X})\bar{\pi}_{\infty}(a\mid\mathbf{X})q^{0}(\mathbf{X})}-1\right\}(\bar{\Lambda}_{\infty}-\bar{\Lambda})(dy\mid a,\mathbf{X})\bigg{]}.$
@@
Following a result in Lemma 1 of Westling et al. [48], $\mathbb{E}\{H_{\bar{S}_{\infty},\bar{G}_{\infty},t,a}(Y,\Delta,\mathbf{X})\mid\mathbf{X}=\mathbf{x}\}$ equals
@@
$+\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{q^{0}_{\infty}(\mathbf{X})}{q^{0}(\mathbf{X})}\bar{S}_{\infty}(t\mid a,\mathbf{X})\frac{\bar{\pi}(a\mid\mathbf{X})}{\bar{\pi}_{\infty}(a\mid\mathbf{X})}$
×$\quad\times\int_{0}^{t}\frac{\bar{S}(y-\mid a,\mathbf{X})\bar{G}(y\mid a,\mathbf{X})}{\bar{S}_{\infty}(y\mid a,\mathbf{X})\bar{G}_{\infty}(y\mid a,\mathbf{X})}(\bar{\Lambda}_{\infty}-\bar{\Lambda})(dy\mid a,\mathbf{X})\bigg{]}.$
Furthermore, by Duhamel equation in Gill and Johansen [22], we have
@@
$\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\bar{S}_{\infty}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{\bar{S}(y-\mid a,\mathbf{X})}{\bar{S}_{\infty}(y\mid a,\mathbf{X})}$
×$\quad\times\left\{\frac{\bar{G}(y\mid a,\mathbf{X})\bar{\pi}(a\mid\mathbf{X})q^{0}_{\infty}(\mathbf{X})}{\bar{G}_{\infty}(y\mid a,\mathbf{X})\bar{\pi}_{\infty}(a\mid\mathbf{X})q^{0}(\mathbf{X})}-1\right\}(\bar{\Lambda}_{\infty}-\bar{\Lambda})(dy\mid a,\mathbf{X})\bigg{]}.$ (5)
D.1.4 Proof of Theorem 2.8
@@
$\widehat{\theta}^{\text{CCOD}}_{n}(t,a)-\theta^{0}(t,a)$ $=\mathbb{P}_{n}[\varphi^{*\text{CCOD}}_{t,a}]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{t,a}\right]$
$\qquad+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}\left[\widehat{\varphi}^{\text{CCOD}}_{t,a}-\theta^{0}(t,a)\right].$
By Conditions D.1 and D.2, the second summand on the right-hand-side is $o_{p}(n^{-1/2})$ by Lemma D.4. By Lemma D.5, $\mathbb{P}[\widehat{\varphi}^{\text{CCOD}}_{t,a}]-\theta^{0}(t,a)$ equals
$\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{\bar{S}(y-\mid a,\mathbf{X})}{\widehat{\bar{S}}_{m}(y\mid a,\mathbf{X})}$
×$\quad\times\left\{\frac{\bar{G}(y\mid a,\mathbf{X})\bar{\pi}(a\mid\mathbf{X})\widehat{q}^{0}_{m}(\mathbf{X})}{\widehat{\bar{G}}_{m}(y\mid a,\mathbf{X})\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})q^{0}(\mathbf{X})}-1\right\}(\widehat{\bar{\Lambda}}_{m}-\bar{\Lambda})(dy\mid a,\mathbf{X})\bigg{]}.$
By Duhamel equation in Gill and Johansen [22], we have that
@@
$=\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})\left\{\frac{\bar{\pi}(a\mid\mathbf{X})}{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})}-1\right\}\int_{0}^{t}\frac{\bar{S}}{\widehat{\bar{S}}_{m}}(du\mid a,\mathbf{X})+\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})$
×$\qquad\times\int_{0}^{t}\frac{\bar{\pi}(a\mid\mathbf{X})}{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})}\left\{\frac{\bar{G}(y\mid a,\mathbf{X})\widehat{q}^{0}_{m}(\mathbf{X})}{\widehat{\bar{G}}_{m}(y\mid a,\mathbf{X})q^{0}(\mathbf{X})}-1\right\}\left(\frac{\bar{S}}{\widehat{\bar{S}}_{m}}-1\right)(du\mid a,\mathbf{X})\bigg{]}$
@@
$\quad+\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{\bar{\pi}(a\mid\mathbf{X})}{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})}\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})$
×$\qquad\underbrace{\quad\times\int_{0}^{t}\left\{\frac{\bar{G}(y\mid a,\mathbf{X})\widehat{q}^{0}_{m}(\mathbf{X})}{\widehat{\bar{G}}_{m}(y\mid a,\mathbf{X})q^{0}(\mathbf{X})}-1\right\}\left(\frac{\bar{S}}{\widehat{\bar{S}}_{m}}-1\right)(du\mid a,\mathbf{X})\bigg{]}}_{I_{2}}.$
@@
$\quad+\mathbb{E}\bigg{[}\frac{\widehat{q}^{0}_{m}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{\bar{\pi}(a\mid\mathbf{X})}{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})}\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})$
×$\qquad\underbrace{\quad\times\int_{0}^{t}\left\{\frac{\bar{G}(y\mid a,\mathbf{X})}{\widehat{\bar{G}}_{m}(y\mid a,\mathbf{X})}-1\right\}\left(\frac{\bar{S}}{\widehat{\bar{S}}_{m}}-1\right)(du\mid a,\mathbf{X})\bigg{]}}_{I_{22}}.$